# Probing Green’s Function Zeros by Co-tunneling through Mott Insulators

### v1

@@ Abstract
In Fig.2 we show the emergent co-tunneling amplitude for the experimental setup in Fig.1(a) between two interacting auxiliary quantum dots, as a function of the bandwidth $D$. We compare their behavior in the case of a Mott insulator and a conventional non-interacting insulator, and plot the two resulting curves (red and grey) as a function of $D$. The numerically obtained curves are to be compared with the two dashed lines, which show the two different behaviors in the non-interacting and Mott limits and can be extracted from Eq. (7). In the non-interacting limit, the behavior of the Green’s function $G$, as defined in Eq.(1), is determined by a single energy scale, the non-interacting bandwidth appearing at the denominator [42, 43]. Hence, $t_{\text{co}}\sim D^{-1}$, and the two solid lines are on top of each other and agree with the non-interacting behavior (negative slope in the logarithmic plot). At around $\frac{D}{U}\approx 1$, i.e. near the Mott transition, the behavior of the interacting system changes: now, the dominant energy scale of the system is the Coulomb interaction $U$, and approximation Eq. (3) holds. $H^{z}$, though renormalized as in Eq. (4), has a $\sim D$ dependence on the bandwidth, and the Mott co-tunneling amplitude indeed agrees with the positive-slope dashed line. The general agreement between the approximated value of $t_{\text{co}}$ and the numerically determined ones is remarkable, confirming the robustness of Eq.(7). Only at small bandwidth $D$ our co-tunneling approximation breaks down ($\gamma\approx D$) and higher order tunneling processes become relevant.
@@
Figure 3: Charge stability diagram (ED, $U=200$ $N=16$, $\gamma=4$,$l_{i}=\delta_{i,2}$,$r_{i}=\delta_{i,3}$, $v=18$, $w=20$,PBC) : (a) Derivative particle number $N_{L}$($N_{R}$) on left(right) dot for varying potentials $\mu_{\text{L}},\mu_{\text{R}}$. In (b): cut through (a) (red line), FWHM is proportional to co-tunneling $t_{\text{co}}$. (c) and (d): charge oscillations between aux. dots (initial localization left dot) same parameter as (a), but OBC, $N=20$, $l_{i}=\delta_{i,1}$,$r_{i}=\delta_{i,2}$ in (c) and $l_{i}=\delta_{i,3}$,$r_{i}=\delta_{i,4}$ in (d). Analytical results (dashed) from Eq.(7).
@@
Our co-tunneling approach to detecting Green’s function zeros can be nicely illustrated with the well established tool of charge-stability diagrams [56, 57, 58, 11, 59], which is commonly used to describe the behavior of quantum dots as a function of control voltages. In Fig.3 (a) we show the charge-stability diagram of our co-tunneling setup, by plotting the derivative of the charge occupation $N_{L/R}$ with respect to the detuning $\delta\mu=\mu_{\text{L}}-\mu_{\text{R}}$ at varying chemical potentials $\mu_{\text{L}},\mu_{\text{R}}$. Due to the finite tunneling amplitude $t_{\text{co}}$, the boundaries of the characteristic checkerboard structure of the uncoupled system ($\gamma=0$) are smeared out and and the localization of a single particle can be tuned smoothly from the left dot $(1,0)$ to the right dot $(0,1)$. Fig.3 (b) shows a diagonal cut through Fig.3 (a) at a fixed average potential $\bar{\mu}=\frac{1}{2}\left(\mu_{\text{L}}+\mu_{\text{R}}\right)$ (red line). The FWHM of the Lorentzian dispersion is directly proportional to the co-tunneling amplitude. An alternative way of probing the co-tunneling amplitude is to measure the frequency of Rabi-charge oscillations between the dots [60, 61, 62, 36], which can be extracted by the time-evolution of a localized state and compared with the theoretical value $\Omega\approx\sqrt{|t_{\text{co}}|^{2}+\omega^{2}_{D}}$, where $\omega_{D}\approx\frac{1}{2}\delta\mu+\mathcal{O}(\frac{\gamma^{2}}{U})$ is a shifted detuning between the two quantum dots, easily controllable via $\mu_{L}$ and $\mu_{R}$. In Fig.3 (c) and (d) we plot the numerically calculated time-dependent charge occupation values of the two quantum dots on resonance $\omega_{D}=0$ (at peak maximum in Fig.3 (b)), when the system is initialized in the completely localized (1,0)-state. The data are in perfect agreement with the predicted oscillation frequency, which can be easily extracted by Eq.(6). In a real experiment, the perturbative limit $\frac{\gamma}{D}\ll\frac{D}{U}\ll 1$, may not be perfectly accessible and is further inherently tied to an small co-tunneling amplitude, making it experimentally challenging. Despite this limitation, macroscopic values of $t_{\text{co}}$ ($\geq 0.5\mu eV$) can still be achieved for a broad range of typical parameters $U,D$ and $\gamma$, within an acceptable quantitative deviation ($\leq 10\%$) from the effective Hamiltonian $\bar{h}$ (see [47] for details). Interestingly, we also find that the striking spin renormalization in Eq.(4) can be probed qualitatively, without requiring fine-tuning of the dots. More specifically, we find a notable amplification of the tunneling amplitude as function of varying boundary conditions as a characteristic signature of the spin-spin correlations (see [47] for further details).
@@ Conclusion
– Starting from a degenerate perturbation theory around the atomic limit, we have derived a simple yet highly accurate approximation for the low-frequency features of a Mott Green’s function, which can be completely described by a dispersion relation for the Green’s function zeros. This situation is reminiscent of a non-interacting system with the important distinction that the effective dispersion of zeros appears in the numerator of the Green’s functions, and is renormalized compared to the underlying free band structure by means of spin correlators, as we have quantified. Our main result is the establishment of a direct link between the dispersion of zeros and the transmission amplitude naturally probed in a co-tunneling experiment, which drastically simplifies the task of observing and characterizing Green’s function zeros. We have provided extensive numerical proof of the correctness and resilience of our approximation, and suggested a feasible experimental setup for the verification of our key findings. Numerical simulations suggest that the fingerprints of the zeros dispersion and renormalization are observable already for small systems: specifically, an ensemble of six quantum dots (2 leads and 4 in the central Hubbard tunneling medium) is already sufficient to quantify an interesting correlation induced renormalization of band parameters.
@@ Appendix A Perturbation theory for the Green’s function
In the following section we derive the GF for the Hubbard model within the strongly correlated limit $\frac{D}{U}\rightarrow 0$ , i.e., deep in the Mott regime. The general Hubbard Hamiltonian is given as:
@@
Deep in the Mott regime, Eq.(8) implies an perturbative expansion [45] of the eigenstates in $\phi=\frac{D}{U}$, where the rescaled bare Hamiltonian $H^{\prime}_{0}=\frac{1}{D}H_{0}$ acts as perturbation. Then, we insert the resulting expanded eigenstates and eigenenergies into the Källen-Lehmann representation and eventually collect all terms up to the second order. Thereby, we will first derive an rescaled GF $G^{\prime}(\phi\frac{\omega}{D})=\frac{D}{\phi}G(\omega)$, where $G^{\prime}$ denotes the GF corresponding to the Hamiltonian $D^{-1}\phi H_{\text{H}}$, afterwards we can easily readout $G(\omega)$.
@@ A.1 Eigenvector structure towards the strongly correlated limit
The Hubbard interaction $H_{\text{int}}$ towards strong correlation $\phi\rightarrow 0$ avoids empty and double occupied sites, i.e., they cost an energy amount of $2D\phi^{-1}=2U$ The eigenenergies will then arrange in well separated energy sectors according the amount of empty and double occupied sides (see Fig.4). We will expand the eigenstates around this strongly correlated limit, hence for house-holding reasons we will index-labelling all the unperturbed eigenstates $\{|E^{0,m}_{\nu,l}\rangle\}$ according to their energy sectors. In the following we denote by $m=N-N_{h}$ the relative particle number, with the total particle number $N=\sum_{j,\sigma}n_{j,\sigma}$ and the particle number at half-filling $N_{h}$. Furthermore, by $l$ we count the empty or double occupied sites additional to the lowest energy state of the corresponding particle sector $m$. In combination, the indices $l$ and $m$ fully determine the energy of each eigenstate in the atomic limit:
@@
Figure 4: Energy levels $E^{0,m}_{\nu,l}=E_{0}+U(|m|+2l)$ for a typical Hubbard system in the Mott regime (atomic limit): all eigenstates can be sorted with respect to the particle-number deviation from half-filling $N=N_{h}+m$ and the amount of empty(douple occupied) sites $l$ additional to the lowest states for a fixed particle-number sector (coloured regions). For each energy there is a huge degeneracy denoted by $\nu$. The energy $E_{0}$ denotes the lowest global energy. In order to get the zero temperature GF the only relevant subspace consist of the global groundstate and the lowest energy sector with $m=|1|,l=0$ (yellow box). Direct Scattering from the groundstate into higher energies with $l>0$ is forbidden within the atomic limit (black arrows).
@@ A.2 Decomposition of the non-interacting Hamiltonian
The bare hopping $H_{0}$ contains various terms which can be resorted with respect to the energy-increase regarding the strongly interacting limit,i.e., their ability to produce empty and double-occupied sites by acting on the groundstate at half-filling $H_{0}|GS_{0}\rangle$.
@@
A.3 First order perturbation theory
As first step we formally expand the perturbed eigenstates(-energies) $|E^{P,m}_{\nu,l}\rangle$($E^{P,m}_{\nu,l}$) of $\frac{\phi}{D}H_{\text{H}}$ in Eq.(8) towards the unperturbed limit, up to the first order in $\phi$. This is essentially standard degenerated perturbation theory [45] and maps the atomic eigenstates(-energies) $|E^{0,m}_{\nu,l}\rangle$($E^{0,m}_{\nu,l}$) smoothly to the perturbed ones.
@@ A.3 First order perturbation theory
Here, we introduce the energy-sector projector $P_{l}^{m}$. It is important to note that the degeneracy within each sector $(m,l)$, denoted by $\nu$, must, in principle, be resolved by applying non-trivial higher-order perturbation theory [45]. However, for the final result, it suffices that such an expansion exists. Moreover, Eq. (13) can be interpreted as an inverse Schrieffer-Wolff transformation [53, 55, 54], which maps the eigenstates of an effective model defined in the atomic limit back to the original basis.
@@
$G^{\prime}_{j,\sigma;j^{\prime},\sigma^{\prime}}(\phi\omega^{\prime})$ $=\sum_{\nu,l}\Bigg{[}\frac{\left<GS^{P}|c_{j,\sigma}|E_{\nu,l}^{P,+}\right>\left<E_{\nu,l}^{P,+}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{P}\right>}{\phi\omega^{\prime}+E_{0}^{P}-E_{\nu,l}^{P,+}}+\frac{\left<E_{\nu,l}^{P,-}|c_{j,\sigma}|GS^{P}\right>\left<GS^{P}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|E_{\nu,l}^{P,-}\right>}{\phi\omega^{\prime}+E_{\nu,l}^{P,-}-E_{0}^{P}}\Bigg{]}$
@@
×$\times\langle E^{0,+}_{\nu,l}|\left(1+\sum_{l^{\prime}\neq l}\frac{\phi}{2(l-l^{\prime})}H^{\prime}_{0}P^{+}_{l}\right)c^{\dagger}_{j^{\prime},\sigma^{\prime}}\left(1-\sum_{k>0}\frac{\phi}{2k}P^{0}_{k}H^{\prime}_{0}\right)|GS^{0}\rangle$
@@
×$\times\langle E^{0,+}_{\nu,l}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}+\phi\left[\sum_{l^{\prime}\neq l}\frac{1}{2(l-l^{\prime})}H^{\prime}_{0}P^{+}_{l^{\prime}}c^{\dagger}_{j^{\prime},\sigma^{\prime}}-c^{\dagger}_{j^{\prime},\sigma^{\prime}}\sum_{k>0}\frac{1}{2k}P^{0}_{k}H^{\prime}_{0}\right]|GS^{0}\rangle$
@@
×$\times\langle E^{0,-}_{\nu,l}|\left(1+\sum_{l^{\prime}\neq l}\frac{\phi}{2(l-l^{\prime})}H^{\prime}_{0}P^{-}_{l}\right)c_{j,\sigma}\left(1-\sum_{k>0}\frac{\phi}{2k}P^{0}_{k}H^{\prime}_{0}\right)|GS^{0}\rangle$
@@
×$\times\langle E^{0,-}_{\nu,l}|c_{j,\sigma}\phi\left[\sum_{l^{\prime}\neq l}\frac{1}{2(l-l^{\prime})}H^{\prime}_{0}P^{-}_{l^{\prime}}c_{j,\sigma}-c_{j,\sigma}\sum_{k>0}\frac{1}{2k}P^{0}_{k}H^{\prime}_{0}\right]|GS^{0}\rangle$
@@
$-\sum_{\nu}A_{j,\sigma;j^{\prime},\sigma^{\prime}}^{0,\nu,0}$ $=-\sum_{\nu}\langle GS^{0}|c_{j,\sigma}|E^{0,+}_{\nu,0}\rangle\langle E^{0,+}_{\nu,0}|c_{j^{\prime},\sigma^{\prime}}^{\dagger}|GS^{0}\rangle=-\langle GS^{0}|c_{j,\sigma}c_{j^{\prime},\sigma^{\prime}}^{\dagger}|GS^{0}\rangle\text{ ,}$ (21)
@@
$=\sum_{k>0}\frac{1}{k}\langle GS^{0}|\left[c_{j,\sigma}c_{j^{\prime},\sigma^{\prime}}^{\dagger},H^{\prime}_{0,k}\right]_{+}|GS^{0}\rangle\text{ ,}$ (22)
@@
$\sum_{\nu}B_{j,\sigma;j^{\prime},\sigma^{\prime}}^{0,\nu,0}$ $=\sum_{\nu}\langle GS^{0}|c_{j^{\prime},\sigma^{\prime}}^{\dagger}|E^{0,-}_{\nu,0}\rangle\langle E^{0,-}_{\nu,0}|c_{j,\sigma}|GS^{0}\rangle=\langle GS^{0}|c_{j^{\prime},\sigma^{\prime}}^{\dagger}c_{j,\sigma}|GS^{0}\rangle\text{ ,}$ (24)
@@
$=-\sum_{k>0}\frac{1}{k}\langle GS^{0}|\left[c_{j^{\prime},\sigma^{\prime}}^{\dagger}c_{j,\sigma},H^{\prime}_{0,k}\right]_{+}|GS^{0}\rangle\text{ ,}$ (25)
@@
$G(\omega)$ $\approx\frac{\phi^{2}}{D^{2}}\left[H^{\text{z}}-\omega\right]+\mathcal{O}(\phi^{3})\text{ ,}$ (33)
@@
A.4 Rewriting the GF approximation Eq.(34) for generic Hubbard Hamiltonians

### v2

@@ Abstract
In Fig.2 we show the emergent co-tunneling amplitude $t_{\text{co}}$ for the experimental setup in Fig.1(a) between two interacting auxiliary quantum dots, as a function of the bandwidth $D$. We compare its behavior in the case of a Mott insulator and a conventional non-interacting insulator, and plot the two resulting curves (red and grey) as a function of $D$. The numerically obtained curves are to be compared with the two dashed lines, which show the two different behaviors in the non-interacting and Mott limits, and can be extracted from Eq. (6). In the non-interacting limit, the behavior of the Green’s function $G$, as defined in Eq. (1), is determined by a single energy scale, the non-interacting bandwidth appearing at the denominator [45, 46]. Hence, $t_{\text{co}}\sim D^{-1}$, and the two solid lines are on top of each other and agree with the non-interacting behavior (negative slope in the logarithmic plot). At around $\frac{D}{U}\approx 1$, i.e., near the Mott transition, the behavior of the interacting system changes: now, the dominant energy scale of the system is the Coulomb interaction $U$, and approximation Eq. (3) holds. $H^{z}$, though renormalized as in Eq. (4), has a $\sim D$ dependence on the bandwidth, and the Mott co-tunneling amplitude indeed agrees with the positive-slope dashed line. The general agreement between the approximated value of $t_{\text{co}}$ and the numerically determined ones is remarkable, confirming the robustness of Eq. (6). Only at small bandwidth $D$, where $\gamma\approx D$, does our co-tunneling approximation break down and higher-order tunneling processes become relevant.
@@
Figure 3: Charge stability diagram (ED, $U=200$ $N=20$, $\gamma=4$$l_{i}=\delta_{i,2}$$r_{i}=\delta_{i,3}$, $v=18$, $w=20$,PBC) : (a) Derivative of the particle number $N_{L}$($N_{R}$) on left(right) dot for varying potentials $\mu_{\text{L}},\mu_{\text{R}}$. In (b): cut through (a) (red line), FWHM is proportional to co-tunneling $t_{\text{co}}$. (c) and (d): charge oscillations between aux. dots (initial localization left dot) same parameter as (a), but OBC, $l_{i}=\delta_{i,1}$$r_{i}=\delta_{i,2}$ in (c) and $l_{i}=\delta_{i,3}$$r_{i}=\delta_{i,4}$ in (d). Analytical results (dashed) from Eq. (6).
@@
Our co-tunneling approach to detecting Green’s function zeros can be nicely illustrated with the well-established tool of charge-stability diagrams [80, 64, 65, 66, 11, 67], which is commonly used to describe the behavior of quantum dots as a function of control voltages. In Fig.3 (a), we show the charge-stability diagram of our co-tunneling setup, by plotting the derivative of the charge occupation $N_{L/R}$ with respect to the detuning $\delta\mu=\mu_{\text{L}}-\mu_{\text{R}}$, at varying chemical potentials $\mu_{\text{L}},\mu_{\text{R}}$. Due to the finite tunneling amplitude $t_{\text{co}}$, the boundaries of the characteristic checkerboard structure of the uncoupled system ($\gamma=0$) are smeared out, and the localization of a single particle can be tuned smoothly from the left dot $(1,0)$ to the right dot $(0,1)$. Fig.3 (b) shows a diagonal cut through Fig.3 (a) at a fixed average potential $\bar{\mu}=\frac{1}{2}\left(\mu_{\text{L}}+\mu_{\text{R}}\right)$ (red line). At zero temperature, the FWHM of the Lorentzian dispersion is directly proportional to the co-tunneling amplitude. An alternative way of probing the co-tunneling amplitude is to measure the frequency of Rabi-charge oscillations between the dots [70, 72, 73, 38]. It can be extracted by the time-evolution of a localized state and compared with the theoretical value $\Omega\approx\sqrt{|t_{\text{co}}|^{2}+\omega^{2}_{D}}$, where $\omega_{D}\approx\frac{1}{2}\delta\mu+\mathcal{O}(\frac{\gamma^{2}}{U})$ is a shifted detuning between the two quantum dots, which is easily controllable via $\mu_{L}$ and $\mu_{R}$. In Fig.3 (c) and (d), we plot the numerically calculated time-dependent charge occupation values of the two quantum dots on resonance $\omega_{D}=0$ (at peak maximum in Fig.3 (b)), when the system is initialized in the completely localized (1,0)-state. The data are in perfect agreement with the predicted oscillation frequency, which can be easily extracted by Eq.(5).
The proposed probing scheme applies to generic Mott insulators, and so far our discussion has been generally applicable. As a particularly promising experimental realization, we now elaborate on a specific Hubbard system synthetically realizing a Mott insulator with coupled quantum dots. In a real experiment on quantum dot arrays, the perturbative limit $\frac{\gamma}{D}\ll\frac{D}{U}\ll 1$, may not be perfectly accessible and is further inherently tied to a small co-tunneling amplitude, making it experimentally challenging. Despite this limitation, macroscopic values of $t_{\text{co}}$ ($\geq 0.5\mu eV$) can still be achieved for a broad range of typical parameters $U,D$ and $\gamma$, within an acceptable quantitative deviation ($\leq 10\%$, see [50] for details). Interestingly, we also find that the striking re-scaling in Eq.(4) by spin-spin correlations can be probed qualitatively, without requiring fine-tuning of the dots. More specifically, we find a notable amplification of the tunneling amplitude as function of varying boundary conditions, a characteristic signature of this renormalization (see [50] for further details).
@@ Conclusion
– Starting from a degenerate perturbation theory around the atomic limit, we have derived a simple yet highly accurate approximation for the low-frequency features of a Mott Green’s function, which can be completely described by a dispersion relation for the Green’s function zeros. This situation is reminiscent of a non-interacting system, with the important distinction that the effective dispersion of zeros appears in the numerator of the Green’s functions, and is renormalized compared to the underlying free band structure by means of spin-spin correlators, as we have quantified. Our main result is the establishment of a direct link between the dispersion of zeros and the transmission amplitude naturally probed in a co-tunneling experiment, which drastically simplifies the task of observing and characterizing Green’s function zeros. Moreover, our approach allows, in principle, a reconstruction of the spectroscopically inaccessible GF determinant from its individual matrix elements. We have provided extensive numerical proof of the correctness and resilience of our approximation, and suggested a feasible experimental setup for the verification of our key findings. Numerical simulations suggest that the fingerprints of the zeros dispersion and renormalization are observable already for small systems: specifically, an ensemble of six quantum dots (2 leads and 4 in the central Hubbard tunneling medium) is already sufficient to quantify an interesting correlation induced renormalization of band parameters.
@@ Appendix A Perturbation theory for the Green’s function
In the following section, we derive the GF for the Hubbard model within the strongly correlated limit $\frac{D}{U}\rightarrow 0$ , i.e., deep in the Mott regime. The general Hubbard Hamiltonian is given as:
@@
Deep in the Mott regime, Eq. (8) implies an perturbative expansion [48] of the eigenstates in $\phi=\frac{D}{U}$, where the rescaled bare Hamiltonian $H^{\prime}_{0}=\frac{1}{D}H_{0}$ acts as perturbation. Then, we insert the resulting expanded eigenstates and eigenenergies into the Källen-Lehmann representation, apply an expansion at low energies and eventually collect all terms up to the second order in $\phi$.
@@ A.1 Eigenvector structure towards the strongly correlated limit
The Hubbard interaction $H_{\text{int}}$ towards strong correlation $\phi\rightarrow 0$ avoids empty and double occupied sites, i.e., they cost an energy amount of $2D\phi^{-1}=2U$ . The eigenenergies will then arrange in well separated energy sectors according the amount of empty and double occupied sides (see Fig.4). We will expand the eigenstates around this strongly correlated limit, hence for house-holding reasons we will index-labelling all the unperturbed eigenstates $\{|E^{0,m}_{\nu,l}\rangle\}$ according to their energy sectors. In the following we denote by $m=N-N_{h}$ the relative particle number, consisting of the total particle number $N=\sum_{j,\sigma}n_{j,\sigma}$ and the particle number at half-filling $N_{h}$. Furthermore, by $l$ we count the empty or double occupied sites additional to the lowest energy state of the corresponding particle sector $m$. In combination, the indices $l$ and $m$ fully determine the energy of each eigenstate in the atomic limit:
@@
Figure 4: Energy levels $E^{0,m}_{\nu,l}=E_{0}+U(|m|+2l)$ for a typical Hubbard system in the Mott regime (atomic limit): all eigenstates can be sorted with respect to the particle-number deviation from half-filling $N=N_{h}+m$, and the amount of empty (douple occupied) sites $l$ additional to the lowest states for a fixed particle-number sector (coloured regions). For each energy there is a huge degeneracy denoted by $\nu$. The energy $E_{0}$ denotes the lowest global energy. In order to get the zero temperature GF, the only relevant subspace consist of the global groundstate and the lowest energy sector with $m=|1|,l=0$ (yellow box). Direct Scattering from the groundstate into higher energies with $l>0$ is forbidden within the atomic limit (black arrows).
@@ A.2 Decomposition of the non-interacting Hamiltonian
The bare hopping $H_{0}$ contains various terms, which can be resorted with respect to the energy-increase regarding the strongly interacting limit,i.e., their ability to produce empty and double-occupied sites by acting on the groundstate at half-filling $H_{0}|GS_{0}\rangle$.
@@
A.3 Green’s function approximation with SU(2)-symmetry
In this section we derive the low-energy GF approximation for SU(2)-symmetric Mott insulators, the more general SU(2)-breaking case is treated in Section A.5. We will first derive an rescaled GF $G^{\prime}(\phi\frac{\omega}{D})=\frac{D}{\phi}G(\omega)$, where $G^{\prime}$ denotes the GF corresponding to the Hamiltonian $D^{-1}\phi H_{\text{H}}$, afterwards we can easily readout $G(\omega)$. We begin by formally expanding the perturbed eigenstates(-energies) $|E^{P,m}_{\nu,l}\rangle$($E^{P,m}_{\nu,l}$) of $\frac{\phi}{D}H_{\text{H}}$ in Eq. (8) towards the unperturbed limit, up to the first order in $\phi$. This is essentially standard degenerated perturbation theory [48] and maps the atomic eigenstates(-energies) $|E^{0,m}_{\nu,l}\rangle$($E^{0,m}_{\nu,l}$) smoothly to the perturbed ones. For the GF approximation the following terms in the eigenstate expansion are sufficient  [^5] In general this expansion is incomplete, however, SU(2)-symmetry in combination with the first order expansion of the GF suppresses lot’s of terms.:
@@ A.3 Green’s function approximation with SU(2)-symmetry
Here, we introduce the energy-sector projector $P_{l}^{m}$. It is important to note that the degeneracy within each sector $(m,l)$, denoted by $\nu$, must, in principle, be resolved by applying non-trivial higher-order perturbation theory [48]. However, for the final result, it suffices that such an expansion exists. Moreover, Eq. (13) can be interpreted as an inverse Schrieffer-Wolff transformation [60, 62, 61], which maps the eigenstates of an effective model, which is defined in the atomic limit, back to the original basis.
@@
$G^{\prime}_{j,\sigma;j^{\prime},\sigma^{\prime}}(\phi\omega^{\prime})$ $=\sum_{\nu,l}\Bigg{[}\frac{\langle GS^{P}|c_{j,\sigma}|E_{\nu,l}^{P,+}\rangle\langle E_{\nu,l}^{P,+}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|GS^{P}\rangle}{\phi\omega^{\prime}+E_{0}^{P}-E_{\nu,l}^{P,+}}+\frac{\langle E_{\nu,l}^{P,-}|c_{j,\sigma}|GS^{P}\rangle\langle GS^{P}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}|E_{\nu,l}^{P,-}\rangle}{\phi\omega^{\prime}+E_{\nu,l}^{P,-}-E_{0}^{P}}\Bigg{]}$
@@
$\times\langle E^{0,+}_{\nu,l}|\left(1+\sum_{l^{\prime}\neq l}\frac{\phi}{2(l-l^{\prime})}H^{\prime}_{0}P^{+}_{l}\right)c^{\dagger}_{j^{\prime},\sigma^{\prime}}\left(1-\sum_{k>0}\frac{\phi}{2k}P^{0}_{k}H^{\prime}_{0}\right)|GS^{0}\rangle$
@@
$\times\langle E^{0,+}_{\nu,l}|c^{\dagger}_{j^{\prime},\sigma^{\prime}}+\phi\left[\sum_{l^{\prime}\neq l}\frac{1}{2(l-l^{\prime})}H^{\prime}_{0}P^{+}_{l^{\prime}}c^{\dagger}_{j^{\prime},\sigma^{\prime}}-c^{\dagger}_{j^{\prime},\sigma^{\prime}}\sum_{k>0}\frac{1}{2k}P^{0}_{k}H^{\prime}_{0}\right]|GS^{0}\rangle$
@@
$\times\langle E^{0,-}_{\nu,l}|\left(1+\sum_{l^{\prime}\neq l}\frac{\phi}{2(l-l^{\prime})}H^{\prime}_{0}P^{-}_{l}\right)c_{j,\sigma}\left(1-\sum_{k>0}\frac{\phi}{2k}P^{0}_{k}H^{\prime}_{0}\right)|GS^{0}\rangle$
@@
$\times\langle E^{0,-}_{\nu,l}|c_{j,\sigma}\phi\left[\sum_{l^{\prime}\neq l}\frac{1}{2(l-l^{\prime})}H^{\prime}_{0}P^{-}_{l^{\prime}}c_{j,\sigma}-c_{j,\sigma}\sum_{k>0}\frac{1}{2k}P^{0}_{k}H^{\prime}_{0}\right]|GS^{0}\rangle$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
A.4 Rewriting the GF-approximation using spin operators